(1+sint ) ( 1 + cost ) = 5 / 4 Find the value of ( 1 - sint ) ( 1 - cost ).

Grade:12

2 Answers

mycroft holmes
8 years ago
My solution disappeared!! Hint: sin t cos t = ½ [(sin t + cost)2-1]. So use substitution sin t +cos t = y, and solve for y. Required expression is 5/4 – 2y
Ajaiy Prabhu
8 years ago
 $2\sin t \cos t + 2 \sin t + 2 \cos t = \frac{1}{2}$, and adding $\sin^2 t + \cos^2t = 1$
$(\sin t + \cos t)^2 + 2(\sin t + \cos t) = \frac{3}{2}$.
Completing the square on the left in the variable $(\sin t + \cos t)$ gives
 $\sin t + \cos t = -1 \pm \sqrt{\frac{5}{2}}$
$\sin t + \cos t = \sqrt{\frac{5}{2}} - 1$
Subtracting twice this from our original equation
$(\sin t - 1)(\cos t - 1) = \sin t \cos t - \sin t - \cos t + 1 = \frac{13}{4} - \sqrt{10}$